OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = 2(x - 3)² - 7?

OpenStudy (anonymous):

You're missing some terms. Assuming you mean y= 2(x-3)(2x-7)?

OpenStudy (anonymous):

no i mean y = 2(x – 3)2 – 7

OpenStudy (anonymous):

Hmm so y=4x-19? It's a straight line, infinite in both directions so the symmetry is at y=-19

terenzreignz (terenzreignz):

I'm positive @BMI229 meant y = 2(x - 3)² - 7

OpenStudy (anonymous):

yes that is what I meant

terenzreignz (terenzreignz):

Okay, @BMI229 You won't always be given the quadratic equations in that form y = a(x - h)² + k But when you are, be thankful, it makes your life easier. What's called the vertex would be located at the point (h , k) The axis of symmetry would be the line x = h

OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = 2(x – 3)2 – 7? x = –6 x = 2 x = 3 x = 7

terenzreignz (terenzreignz):

mhmm... so what is the h-value in this equation y = 2(x - 3)² - 7 ?